# Pushing Large Language Models to the 6G Edge: Vision, Challenges, and Opportunities

### v1

@@ Abstract
Large language models (LLMs), which have shown remarkable capabilities, are revolutionizing AI development and potentially shaping our future. However, given their multimodality, the status quo cloud-based deployment faces some critical challenges: 1) long response time; 2) high bandwidth costs; and 3) the violation of data privacy. 6G mobile edge computing (MEC) systems may resolve these pressing issues. In this article, we explore the potential of deploying LLMs at the 6G edge. We start by introducing killer applications powered by multimodal LLMs, including robotics and healthcare, to highlight the need for deploying LLMs in the vicinity of end users. Then, we identify the critical challenges for LLM deployment at the edge and envision the 6G MEC architecture for LLMs. Furthermore, we delve into two design aspects, i.e., edge training and edge inference for LLMs. In both aspects, considering the inherent resource limitations at the edge, we discuss various cutting-edge techniques, including split learning/inference, parameter-efficient fine-tuning, quantization, and parameter-sharing inference, to facilitate the efficient deployment of LLMs. This article serves as a position paper for thoroughly identifying the motivation, challenges, and pathway for empowering LLMs at the 6G edge.
@@ I Introduction
The rise of large language models (LLMs), fueled by the success of transformers, has sparked significant interest in the AI community and the whole world. Nowadays, major players in the AI industry are vying to develop their own LLMs, with notable examples including OpenAI’s GPT-3, Google’s PALM, and Meta’s LLaMA. These LLMs, trained on extensive and diverse datasets from the Internet, exhibit unparalleled generalization capabilities when the model size substantially increases – a phenomenon known as “emergence”. For instance, benefiting from its staggering model size, GPT-3 could successfully multiply numbers, even though they were not explicitly trained to do so [1]. Due to their exceptional capabilities, the models can be directly applied or easily adapted (e.g., fine-tuned or instruction tuning) to numerous downstream/unseen tasks, unlocking unprecedented potential in various applications, such as Chatbot, content generation, healthcare, and robotics.
@@
Unfortunately, the existing LLM products purely rely on cloud computing, which suffers from excessive latency, high bandwidth cost, and severe privacy concerns, significantly hampering the effective application of LLMs. First of all, it is infeasible to support fast model inference for real-time applications (e.g., LLM-empowered robotics control/navigation/exploration [2]) based on cloud computing due to the need for timely response. Second, the emergence of multimodal LLMs requires input/output of not only texts, but also images, videos, audio, and other sensory data [3]. Centralizing the massive data for either training or inference will consume significant backhaul/backbone network bandwidth and put great pressure on the central cloud. At last, LLM training or inference raises severe privacy concerns, particularly considering that the data could involve highly sensitive data, such as medical data or human activities including audio instructions and gestures at home. As a result, there is an urgent need to leverage mobile edge computing (MEC) and distributed learning to adapt and deploy LLMs on or in closer proximity to data sources while preserving data ownership of end users.
@@
As we are progressing towards the early standardization of 6G, it is widely recognized that 6G will evolve into a mobile network supporting in-network and distributed AI at the edge [4]. However, considering the intensive computing workload of LLMs, is it even feasible to run such large models at the 6G edge? Thanks to various cutting-edge AI technologies, the answer is yes. We argue that split machine learning (including both split learning and inference) has the potential to fulfill computing needs by partitioning the intensive workload over distributed edge devices/servers. In addition, several mature AI techniques, such as parameter-efficient fine-tuning and model quantization, can substantially reduce communication, computation, and memory requirements for model training (fine-tuning) and inference. For instance, by combining parameter-efficient training and quantization, quantized low-rank adapters [5] (QLoRA) can successfully fine-tune a state-of-the-art LLM on a single consumer GPU for a downstream dataset within 24 hours or even less (depending on the model size), while achieving performance comparable to state-of-the-art LLMs such as GPT-3 [5]. Many ongoing industrial efforts are made to realize on-device LLMs of compact versions. All these facts demonstrate the viability and great potential of adapting or deploying LLMs at the mobile edge.
@@
The convergence of LLM deployment and 6G MEC systems will be an exciting research area. Some prior papers, such as [6, 7], mainly discuss how to leverage LLMs to optimize wireless networks or MEC (i.e., LLMs for networks). To our best knowledge, this is the first article that focuses on how to leverage 6G MEC to support LLM training and inference (i.e., networks for LLMs). Particularly, we will comprehensively examine the promising AI technologies to realize such implementations and elaborate on the integrated communication-computing design for LLMs in mobile edge networks.
@@ II Killer Applications
Healthcare is widely recognized as a crucial application for LLMs. Google’s Med-PaLM 2, for example, is an LLM fine-tuned on medical datasets, capable of delivering high-quality answers to medical inquiries [8]. Med-PaLM 2 surpasses the pass mark on the US Medical License Exam (USMLE) and obtained 86.5% accuracy. Indeed, with multimodal inputs and outputs, LLMs can function as AI medical generalists, offering a variety of healthcare services to users, ranging from chatbots to diagnosis to early warnings [3]. It is thrilling to envision a future where everyone can have their own personal health AI expert to constantly monitor their well-being and provide timely advice. Nevertheless, the massive multimodal data transmissions may pose significant challenges to the cloud-based healthcare LLM deployment. More importantly, cloud-based centralized training or inference faces substantial challenges in collecting data in the medical domain owing to privacy concerns and data regulations, which necessitates privacy-preserving distributed learning, such as federated and split learning, to train/deploy models at the edge.
@@ V Efficient Large Model Training at the Edge
With LLMs pre-trained by the cloud, edge training can fine-tune them to new environments and personalize towards individual needs with local data samples. Despite of the significantly reduced workload compared with training from scratch, fine-tuning LLMs still presents significant challenges to the 6G edge. In this section, we discuss three training techniques to enable fine-tuning of LLMs in 6G edge networks.
V-A Parameter-efficient Fine-tuning
@@
To adapt well-trained models to new tasks/environments, conventional full-parameter fine-tuning (i.e., updating all parameters) for LLMs is computationally expensive for an MEC server. Besides, for distributed or federated learning with model aggregation, full-parameter fine-tuning also incurs considerable communication costs for model transfer. To mitigate these issues, the network operator can employ parameter-efficient fine-tuning. Specifically, by updating only a small proportion of parameters, an LLM can be efficiently adapted to new tasks/environments, thereby significantly decreasing the training/communication overhead while preventing overfitting. As shown in Fig. 2, there are several representative parameter-efficient fine-tuning approaches for LLMs, including adapter tuning, prompt tuning, and Low-rank adaptation (LoRA). Adapter tuning involves inserting well-designed adapter modules between layers for training, while prompt tuning adds tunable prefix tokens. LoRA decomposes attention weight updates into low-rank matrices for updating. The common principle of these methods is to merely train a small amount of parameters, usually less than 1% of the original parameters, thereby dramatically reducing the number of trainable parameters. For instance, applying LoRA to GPT-3 could lead to a remarkable reduction in trainable parameters, falling from 175.2 billion to 37.7 million – a nearly 4600-fold reduction. As shown in Fig. 3, by integrating LoRA and federated split learning (which will be introduced in the next subsection), we observe that a higher freezing ratio potentially compromises model performance since less trainable parameters can be tuned. However, it significantly decreases computing and communication latency at the mobile edge. Consequently, it is essential to choose an optimal freezing ratio to strike the optimal trade-off between accuracy and latency under communication and computing constraints at the edge.
@@
Figure 3: The performance for training latency and bilingual evaluation understudy (BLEU, a metric for evaluating the machine translations against the human translations) of federated split learning versus the freezing ratio, where LoRA is employed to fine-tune GPT-2 medium on WebText dataset. An edge server and 20202020 clients are considered. Computing capabilities of clients and the edge server are set to 3.56 and 35.6 (peak performance of one NVIDIA RTX 3090) TFLOPS, uplink and downlink rates are 70Mbps and 300Mbps, and the number of tokens utilized for training is 264M.
V-B Split Edge Learning
While parameter-efficient fine-tuning can greatly help, the latency might still be excessive if merely utilizing a single edge server. Besides, directly uploading the raw data from edge devices to an edge server for training raises significant privacy concerns. To address these issues, split learning (SL) emerges as a complementary solution.
The original goal of SL is mainly to mitigate privacy leakage. Despite of data privacy preservation, on-device training, such as federated learning (FL), is impractical for LLM training due to the limited resources on edge devices. Instead, SL partitions a model into two sub-models and places them on clients and a server for collaborative training, as illustrated in Fig. 1. In such a case, an edge device trains several early layers, which not only prevents raw data sharing with the server but also makes local training much more affordable. While the vanilla SL trains models between a server and clients in a sequential manner, later variants of SL, including parallel split learning (PSL) and split federated learning (SFL), parallelize the framework by enabling multiple devices to train a model with a server simultaneously, thereby further accelerating the process. A detailed review of the integration of SL and MEC systems, named split edge learning (SEL), can be found in [11].
Nevertheless, SL with two sub-models may still be unable to effectively support LLM training. By extending to multi-hop SL, multiple edge servers can work collaboratively to further partition the heavy training workload of an LLM, as illustrated in Fig. 4. Specifically, several edge servers can form a mesh of computing network for performing multi-hop SL, thereby sharing the heavy computing workload among a set of edge servers. Fig. 5 demonstrates the total training time (including both computing and communication latency) for achieving a target accuracy. As can be seen, although leveraging more edge servers takes full advantage of distributed computing resources, it also increases communication overhead due to the additional smashed data transmission between the servers. For this reason, three edge servers perform better than four edge servers in terms of end-to-end latency in Fig. 5. Consequently, a 6G network operator needs to carefully design how many sub-models the original model should be partitioned into, which layers to split, and where to place the sub-models under the resource-constrained mobile edge. The judicious model splitting and placement play a crucial role in improving the efficiency of multi-hop SL for LLMs.
V-C Quantized Training
@@
Apart from parameter-efficient fine-tuning and SL, quantized training provides another promising solution for LLM training by reducing communications, training, and memory requirements. First of all, under distributed learning with model synchronization requirements, quantized training techniques such as QSGD [12] can significantly alleviate the communication burden for model training at the mobile edge because of the compressed payload. Second, low-precision computation enabled by quantized training can accelerate training speed and reduce energy consumption, making it extremely useful for training LLMs on resource-constrained edge devices/servers. For example, fully quantized training (FQT) [13] replaces the original full-precision computations with low-precision computations by quantizing weights, activations, and gradients. Note that specialized hardware is important for realizing the full benefits of low-precision operations. At last, quantized training enables memory-efficiency model training on edge devices. Combining quantization techniques with parameter-efficient methods can significantly reduce memory footprint during fine-tuning pre-trained large models at the edge. For instance, QLoRA [5] quantizes the model to 4 bits and fine-tunes it using 16-bit low-rank adapters. As a result, it allows fine-tuning a 65B parameter LLM on a 48GB GPU within 24 hours, while achieving comparable performance (e.g., 99.3%) to ChatGPT on the evaluated dataset.
@@
The aforementioned methods focus on model quantization itself, particularly on a single device. Under 6G MEC systems, multiple users often train a model collaboratively with the assistance of an edge server, particularly based on SL or SFL. In such a case, model quantization for different users directly influences the optimal allocation of communication and computing resources at the mobile edge. Hence, a joint design of quantization ratio, bandwidth allocation, and computing resource management plays an essential role in striking the balance between accuracy and latency, thereby boosting the training performance at the resource-constrained network edge. Along this line, an accurate model that captures the impact of the quantization ratio on model convergence is needed. Additionally, the time-varying network conditions during the entire training period must be considered when determining the best quantization ratios for training LLMs.
@@ VI Efficient Large Model Inference at the Edge
VI-A Edge Split Inference
Analogous to SL which aims at model training, split inference is a model inference technique that offloads the computing workload from edge devices to a server via layer-wise model partitioning. Moreover, by uploading smashed data rather than raw data, split inference can lower communication overhead if the size of split layer output is smaller than the size of raw data, which is beneficial for multimodal LLMs with high-definition images or videos as the input data.
In the era of LLMs, a 6G network operator can employ multi-hop split inference to fully harness the distributed computing resources. Supposing there is a sentence or a batch of images for inference, an edge server may fail to calculate the output in a timely fashion due to limited computing resources or lack of memory space. Multi-hop split inference can harness the computing resources of multiple edge servers to execute LLM inference, which improves the service response and shares memory usage among the edge servers. Similar to Section V-B, given a specific network topology with bandwidth constraints, the joint problem of model splitting and placement is also crucial for multi-hop split inference to achieve the best orchestration of network-wide computing resources.
VI-B Quantized Edge Inference
@@
Apart from split inference, another feasible solution for fast LLM inference is model compression. There are various model compression approaches for efficient model inference (e.g., model quantization and pruning), and we only focus on model quantization to illustrate the key issues associated with the network edge. There exist two mainstream quantization techniques for efficient model inference, i.e., quantization-aware training (QAT) and post-training quantization (PTQ) [14]. The former emulates quantization errors during the training process, leading to benefits during the inference stage of low-precision models, whereas the latter directly downsamples the parameters of a well-trained model to the low-bit version. In comparison, QAT generally results in better model performance while PTQ is much less computationally expensive as no retraining is required. It is noted that QAT aims to enhance the accuracy of lower-precision models by emulating inference-time quantization errors during training, which fundamentally differs from quantized training in Section V-C with the goal of reducing training costs.
@@
To deploy LLMs with various precision levels for edge inference, PTQ would be a better option due to its efficiency. In general, QAT can be too costly as it requires re-training/fine-tuning of full-precision LLMs for low-bit deployment. Note that a cloud center cannot always provide different bit versions of models for downloading because the mobile edge could have its own context-aware LLMs adapted to local environments. Instead, the mobile edge can utilize PTQ, allowing the conversion of a high-precision LLM to its low-precision counterparts without any re-training. Upon receiving an inference request, the edge system can customize the bit precision for inference according to network resource availability and users’ QoS requirements in terms of inference latency and accuracy, and then transmit the corresponding models with the desired precision to the devices/servers for inference. In general, high-precision (sub) models can be deployed on devices/servers with powerful communication/computing capabilities, otherwise low-precision (sub) models can be deployed. Additionally, weight quantization is often a preferable option for LLMs, because weight quantization generally leads to smaller performance degradation than activation quantization when the model size is large [14].
VI-C Parameter-sharing Edge Inference
GPU memory will be one of the most precious resources at the network edge. For model inference, model layers need to be loaded into the memory, and the required memory increases with the model size. Consequently, performing inference for LLMs requires a significant amount of running memory. For example, a single inference task for FP16 GPT 6.7B with 512 input sequences and 32 output sequences approximately requires 41.84 GB of running memory when the batch size is set as 64 [9]. The powerful consumer GPU, NVIDIA RTX 4090, only has 24 GB of memory space. As the number of users increases and the sequence length grows, the running memory required will become substantial.
As mentioned in Section IV-B, parameter sharing is a common feature for LLMs. When multiple models share the same parameters/blocks, only one copy of the shared parameters needs to be loaded into GPU memory, thereby significantly reducing the number of swaps and the cost per swap for inference and the required memory space [15]. For instance, by considering seven models (YOLO, Faster RCNN, ResNet, VGG, SSD, Inception, and Mobilenet) with parameter sharing for object classification and detection, the running memory decreases by up to 86.4%, saving 9.9 GB memory [15]. These results can be easily extended to LLMs with shared parameters due to the popularity of parameter-efficient fine-tuning in LLM training.
While parameter sharing saves memory space, it suffers from accuracy degradation since models would have fewer trainable parameters for specific tasks. In general, inference accuracy declines as the number of shared layers rises [15]. This necessitates the selection of appropriate sharing ratios to strike the balance between inference accuracy and memory usage.

### v2

@@ Abstract
Large language models (LLMs), which have shown remarkable capabilities, are revolutionizing AI development and potentially shaping our future. However, given their multimodality, the status quo cloud-based deployment faces some critical challenges: 1) long response time; 2) high bandwidth costs; and 3) the violation of data privacy. 6G mobile edge computing (MEC) systems may resolve these pressing issues. In this article, we explore the potential of deploying LLMs at the 6G edge. We start by introducing killer applications powered by multimodal LLMs, including robotics and healthcare, to highlight the need for deploying LLMs in the vicinity of end users. Then, we identify the critical challenges for LLM deployment at the edge and envision the 6G MEC architecture for LLMs. Furthermore, we delve into two design aspects, i.e., edge training and edge inference for LLMs. In both aspects, considering the inherent resource limitations at the edge, we discuss various cutting-edge techniques, including split learning/inference, parameter-efficient fine-tuning, quantization, and parameter-sharing inference, to facilitate the efficient deployment of LLMs. This article serves as a position paper for thoroughly identifying the motivation, challenges, and pathway for enabling LLMs at the 6G edge.
@@ I Introduction
The rise of large language models (LLMs), fueled by the success of transformers, has sparked significant interest in the AI community and the whole world. Nowadays, major players in the AI industry are vying to develop their own LLMs. Notable examples include OpenAI’s GPT-3, Google’s PALM, and Meta’s LLaMA. These LLMs, trained on extensive and diverse datasets from the Internet, exhibit generalization capabilities when the model size substantially increases – a phenomenon known as “emergence”. For instance, benefiting from its staggering model size, GPT-3 could successfully multiply numbers, even though they were not explicitly trained to do so [1]. Due to their exceptional capabilities, the models can be directly applied or easily adapted (e.g., fine-tuned or instruction tuning) to numerous downstream/unseen tasks, thereby unlocking unprecedented potential in various applications, such as Chatbot, content generation, healthcare, and robotics.
@@
Unfortunately, the existing LLM products purely rely on cloud computing, which suffers from excessive latency, high bandwidth cost, and severe privacy concerns. First of all, it is infeasible to support fast model inference for real-time applications (e.g., LLM-empowered robotics control/navigation/exploration [2]) based on cloud computing due to the need for timely response. Second, the emergence of multimodal LLMs requires input/output of not only texts, but also images, videos, audio, and other sensory data [3]. Centralizing the massive data for either training or inference will consume significant backhaul/backbone network bandwidth and put great pressure on the central cloud, which is not scalable. At last, LLM training or inference raises severe privacy concerns, particularly considering that the data could involve highly sensitive data, such as medical data or human activities including audio instructions and gestures at home. As a result, there is an urgent need to leverage mobile edge computing (MEC) to finetune and deploy LLMs on or in closer proximity to data sources while preserving data ownership of end users.
@@
As we are progressing towards the early standardization of 6G, it is widely recognized that 6G will evolve into a mobile network supporting in-network and distributed AI at the edge [4]. However, considering the intensive computing workload of LLMs, is it even feasible to run such large models at the 6G edge? Thanks to various cutting-edge AI technologies, the answer is yes. We argue that split machine learning (including both split learning and inference) has the potential to fulfill computing needs by partitioning the intensive workload over distributed edge devices/servers. In addition, several mature AI techniques, such as parameter-efficient fine-tuning and model quantization, can substantially reduce communication, computation, and memory requirements for model training (fine-tuning) and inference. For instance, by combining parameter-efficient training and quantization, quantized low-rank adapters (QLoRA) [5] can successfully fine-tune a 65B LLM on a single consumer GPU for a downstream dataset within 24 hours, while achieving performance comparable to state-of-the-art LLMs such as GPT-3 [5]. Moreover, many ongoing industrial efforts are made to deploy sub-10 billion parameter LLMs on phones. All these facts demonstrate the viability and great potential of adapting or deploying LLMs at the mobile edge.
@@
The convergence of LLM deployment and 6G MEC systems will be an exciting research area. Some prior papers, such as [6, 7], mainly discuss how to leverage LLMs to optimize wireless networks or MEC (i.e., LLMs for networks). On the other hand, this article focuses on how to leverage 6G MEC to support LLM training and inference (i.e., networks for LLMs). Particularly, we will elaborate on the killer MEC applications, the tailored MEC architecture, and the integrated communication-computing technologies for deploying LLMs in mobile edge networks.
@@ II Killer Applications
Healthcare is widely recognized as a crucial application for LLMs. Compared to traditional AI models, LLMs exhibit exceptional generalization capabilities, fostering more effective interactions with patients, caregivers, and medical professionals. Google’s Med-PaLM 2, for example, is an LLM fine-tuned on medical datasets, capable of delivering high-quality answers to medical inquiries. Med-PaLM 2 surpasses the pass mark on the US Medical License Exam (USMLE) and obtained 86.5% accuracy. Indeed, with multimodal inputs and outputs, LLMs can function as AI medical generalists, offering a variety of healthcare services to users, ranging from chatbots to diagnosis to early warnings [3]. It is thrilling to envision a future where everyone can have their own personal health AI expert to constantly monitor their well-being and provide timely advice. Nevertheless, the massive multimodal data transmissions may pose significant challenges to the cloud-based healthcare LLM deployment. More importantly, cloud-based centralized training or inference faces substantial challenges in collecting data in the medical domain owing to privacy concerns and data regulations, which necessitates privacy-preserving distributed learning, such as federated and split learning, to train/deploy models at the edge.
@@ V Efficient Large Model Training at the Edge
With LLMs pre-trained by the cloud, edge training can fine-tune them to new environments and personalize towards individual needs. Compared with training from scratch, fine-tuning LLMs requires significantly less training costs. However, it still presents significant challenges to the 6G edge. In this section, we discuss three training techniques to enable fine-tuning of LLMs in 6G edge networks.
V-A On-device (On-server) Training
@@ V-A1 Parameter-efficient fine-tuning
To fine-tune LLMs on edge devices/servers, conventional full-parameter fine-tuning (i.e., updating all parameters) proves to be computationally expensive. Besides, for distributed or federated learning, full-parameter fine-tuning also incurs considerable communication costs associated with model aggregation. To address these concerns, network operators can implement parameter-efficient fine-tuning techniques. Specifically, by updating only a minor fraction of parameters, an LLM can be effectively adapted to new tasks or environments, significantly reducing training and communication overhead while simultaneously mitigating overfitting. As shown in Fig. 2, there are several representative parameter-efficient fine-tuning approaches for LLMs, including adapter tuning, prompt tuning, and Low-rank adaptation (LoRA). Adapter tuning involves inserting well-designed adapter modules between layers for training, while prompt tuning adds tunable prefix tokens. LoRA decomposes attention weight updates into low-rank matrices for updating. The shared principle among these methods is to train a small quantity of parameters, typically less than 1% of the original parameters, which dramatically decreases the number of trainable parameters.
@@
Figure 3: The performance for training latency and bilingual evaluation understudy (BLEU, a metric for evaluating the machine translations against the human translations) of federated split learning versus the freezing ratio, where LoRA is employed to fine-tune GPT-2 medium on WebText dataset. An edge server and $2020$ clients are considered. Computing capabilities of clients and the edge server are set to 3.56 and 35.6 (peak performance of one NVIDIA RTX 3090) TFLOPS, uplink and downlink rates are 70Mbps and 300Mbps, and the number of tokens utilized for training is 264M.
V-A2 Quantized training
@@
Apart from parameter-efficient fine-tuning, quantized training provides another promising solution for LLM training by reducing communications, training, and memory requirements. First of all, under distributed learning with model synchronization requirements, quantized training techniques such as QSGD [10] can significantly alleviate the communication burden benefiting from the compressed models/gradients. Second, low-precision computation can accelerate training speed and reduce energy consumption. For example, fully quantized training (FQT) [11] replaces the original full-precision computations with low-precision computations by quantizing weights, activations, and gradients. At last, quantized training also enables memory-efficiency model training on edge devices by reducing bitwidths.
A recent trend in quantized training for on-device LLM training is the integration of quantization and parameter-efficient fine-tuning to significantly reduce memory footprint. For instance, QLoRA [5] can quantize a model to a 4-bit model and fine-tune it using 16-bit low-rank adapters. As a result, it allows fine-tuning a 65 billion parameters LLM on a 48GB GPU within 24 hours, while achieving comparable performance (e.g., 99.3%) to ChatGPT on the evaluated dataset. This demonstrates the feasibility of fine-tuning LLMs on an edge server or even on an edge device.
V-B Device-server Co-training
While the aforementioned techniques can greatly help, on-device LLM fine-tuning can still be extremely resource-intensive for most edge devices. On the other hand, directly uploading the raw data from edge devices to an edge server for training raises significant privacy concerns. To address these concerns, device-server co-training can be a promising LLM training paradigm in mobile edge computing systems.
To facilitate large model training while preserving the data privacy of edge devices, split learning (SL) can come to rescue. SL partitions a model into two sub-models and places them on clients and a server for collaborative training, as illustrated in Fig. 1. This approach not only renders local training considerably more affordable, but also significantly enhances data privacy, given that only cut-layer smashed data (features) are transmitted to the server. While the vanilla SL trains models between a server and clients sequentially, later variants of SL, including parallel split learning (PSL) and split federated learning (SFL), parallelize the framework by enabling multiple devices to train a model with a server simultaneously, thereby further accelerating the process. A detailed review of the integration of SL and MEC systems, named split edge learning (SEL), can be found in [12]. As shown in Fig. 3, by integrating LoRA (the aforementioned parameter-efficient fine-tuning technique) and SFL, we observe model training can be conducted in a reasonable time even if only an NVIDIA RTX 3090 is considered on the edge server. Moreover, a higher freezing ratio potentially compromises model performance while significantly decreasing communication-computing latency at the mobile edge. Consequently, it is essential to choose an optimal freezing ratio to strike the optimal trade-off between accuracy and latency under communication and computing constraints, which can be a significant future research direction.
In general, to balance the communication and computing workload, the optimal model splitting strategies can be formulated as integer optimization problems which are normally challenging to tackle. Fortunately, for LLMs built on transformers, the output sizes of transformer blocks are typically the same (unlike CNNs with varied output sizes for each layer), allowing us to design efficient and optimal model splitting algorithms for LLMs benefiting from the special model architecture. Besides, when wireless communication is the bottleneck, SL for LLMs can be implemented by simply placing the encoders on edge devices and the decoders on the server to reduce communication latency. This is because only tokens, which are usually very small in size, need to be exchanged in this case.
@@ V-B Device-server Co-training
SL with two sub-models may still be unable to support LLM training effectively. By extending to multi-hop SL, multiple edge servers can work collaboratively to further partition the heavy training workload of an LLM, as illustrated in Fig. 4. Specifically, several edge servers can form a mesh of computing network for performing multi-hop SL, thereby sharing the heavy computing workload among a set of edge servers. Fig. 5 demonstrates the total training time (including both computing and communication latency) for achieving a target accuracy. As can be seen, although leveraging more edge servers takes full advantage of distributed computing resources, it also increases communication overhead due to the additional smashed data transmission between the servers. For this reason, three edge servers perform better than four edge servers in terms of end-to-end latency in Fig. 5. Consequently, a 6G network operator needs to carefully design how many sub-models the original model should be partitioned into, which layers to split, and where to place the sub-models under the resource-constrained mobile edge. The judicious model splitting and placement play a crucial role in improving the efficiency of multi-hop SL for LLMs, which is worth further exploration.
@@ VI Efficient Large Model Inference at the Edge
VI-A On-device (On-server) Inference
VI-A1 Quantized edge inference
@@
A feasible solution for fast LLM inference is model quantization. There exist two mainstream quantization techniques for efficient model inference, i.e., quantization-aware training (QAT) and post-training quantization (PTQ). The former emulates quantization errors during the training process, leading to benefits during the inference stage of low-precision models, whereas the latter directly downsamples the parameters of a well-trained model to the low-bit version. In comparison, QAT generally results in better model performance while PTQ is much less computationally expensive as no retraining is required. It is noted that QAT aims to enhance the accuracy of lower-precision models by emulating inference-time quantization errors during training, which fundamentally differs from quantized training in Section V-A with the goal of reducing training costs.
@@
For resource-limited edge devices/servers, PTQ offers a better option than QAT in terms of efficiency. QAT involves re-training/fine-tuning full-precision LLMs for low-bit deployment, which might be too costly if the fine-tuning should be done on the mobile edge. Instead, the mobile edge can utilize PTQ, allowing the conversion of a high-precision LLM to its low-precision counterparts without re-training. This approach also reduces edge caching costs since only a high-precision LLM needs to be stored on edge servers/devices. Upon receiving an inference request, the edge system can customize the bit precision for inference according to network resource availability and users’ QoS requirements in terms of inference latency and accuracy, and then transmit the corresponding downsampled models with the desired precision to the devices/servers for on-device inference.
Apart from the traditional quantization approaches, quantization methods tailored for LLMs on edge devices/servers have also been developed. A recent trend in LLMs is to deploy mixture-of-expert (MoE) LLMs, which have a set of “experts” to be selected at runtime via small-sized non-experts weights, thereby scaling the model parameter size with sublinearly increased computing complexity. Since a substantial portion of the computational load is carried by weights “non-experts”, edge devices/servers can hold only non-expert weights in memory and keep experts in its disk, thereby switching experts between memory and disk as they are required [13]. In such a case, the quantization hence can be customized for different experts according to their sensitivity to quantization errors and the memory budget, thereby reducing swapping time.
VI-A2 Parameter-sharing edge inference
An edge device/server may run multiple LLM-empowered applications at the same time, which requires a significant amount of running memory. For example, a single inference task for FP16 GPT 6.7B with 512 input sequences and 32 output sequences approximately requires 41.84 GB of running memory when the batch size is set as 64 [8]. The powerful consumer GPU, NVIDIA RTX 4090, only has 24 GB of memory space. As the number of running applications increases and the sequence length grows, the running memory requirement will be a major bottleneck for edge devices/servers. Thankfully, as discussed in Section IV-B, parameter sharing is a prevalent feature among LLMs. By exploiting this property, when multiple models consist of the same parameters or blocks, we can load only a single copy of the shared parameters into the GPU memory, thereby substantially reducing memory costs and data swapping time [14]. To this aim, task scheduling for parameter-sharing edge inference can be developed to fully exploit the parameter shareability of models/tasks.
VI-B Device-server Co-inference
To further reduce the inference time and memory costs, device-server co-inference can be deployed at the mobile edge. Split inference is a model inference technique that offloads the computing workload from edge devices to a server via layer-wise model partitioning. In addition to good privacy preservation, split inference can reduce communication overhead if the size of the cut-layer features is smaller than the size of the raw data. This is particularly advantageous for multimodal LLMs that use high-definition images or videos as input data. To further reduce service time, a 6G network operator can employ edge devices and multiple edge servers for collaborative inference based on multi-hop split inference to fully harness the distributed computing resources. Similar to Section V-B, given a specific network topology with bandwidth constraints, the joint problem of model splitting and placement is also crucial for multi-hop split inference to achieve the best orchestration of network-wide computing resources.
Device-server co-inference can also be executed based on the state-of-the-art speculative decoding process of LLMs to significantly reduce inference latency [15]. Specifically, large language models typically generate tokens in an autoregressive manner, which requires running the model K times to produce K tokens, resulting in significant latency. Speculative decoding addresses this issue by using a smaller model to generate multiple tokens, which are then verified and corrected by the larger model in parallel. This method is much faster than the standard autoregressive generation in LLMs, as it leverages the small model’s quick token generation and the large model’s parallel verification and correction capabilities. In the mobile edge, small models can be deployed on edge devices to execute tasks with low latency and low memory/storage costs, after which the generated tokens are uploaded to an edge server for verification in parallel. Since only tokens are transferred (e.g., text), the wireless transmission usually incurs a very small delay. This device-server co-inference paradigm not only accelerates the overall process by facilitating parallel verification through speculative decoding itself but also produces local prediction results on devices via small models for timely response. This allows edge devices to utilize the low-latency results before receiving more accurate results from an edge server.